OpenStudy (anonymous):

rationalize the denominator. square root of 6 over the square root of 12?

OpenStudy (anonymous):

When you rationalize the denominator that means get rid of any square roots that are in it. So if you have \[\sqrt{6}/\sqrt{2}\] you need to get rid of the square root sign in the denominator. How? well \[\sqrt{2}\sqrt{2}\] so if you multiply top and bottom of the fraction √6/√2 by √2 (you are basically multiplying by 1) you rationalize the denominator. So